OpenStudy (anonymous):

Write the rate as a unit rate. James Saltwood can paint 98 doors in 7 hours, while Scott Kepler can paint 38 doors in 3 hours. Who can paint doors faster, James or Scott?

OpenStudy (amistre64):

98/7 38/3 ---- compared to ---- 7 3

OpenStudy (anonymous):

98/7 = doors/ per hour 38/3 = doors/per hour

OpenStudy (anonymous):

A. Scott B. James

OpenStudy (anonymous):

so scott is faster?

OpenStudy (amistre64):

10*3 = 30 4*7 = 28 to guestimate

OpenStudy (amistre64):

id pick james maybe

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Scott : 98/ 7 = 14 doors/hr James: 38/ 3 = 12.7 doors/hr => Scott faster

OpenStudy (anonymous):

Sorry, messed up with the names: James : 98/ 7 = 14 doors/hr Scott : 38/ 3 = 12.7 doors/hr => Jame's faster